Differentiate with respect to the variable 


y=x3 - x2 + x + 5. dy/dx


y=x3 + 7x2 - 4x + 8 dy/dx


s=17t3 - 5t2 - 5t - 3 ds/dt




1
Expert's answer
2021-04-15T07:21:27-0400

Solution:

(a):

"y=x^3 - x^2 + x + 5"

On differentiating both sides w.r.t "x" ,

"\frac{dy}{dx}=3x^2-2x+1+0" [Using "\frac{d}{dx}(x^n)=n.x^{n-1}" ]

"\Rightarrow \frac{dy}{dx}=3x^2-2x+1"

(b):

"y=x^3 + 7x^2 - 4x + 8"

On differentiating both sides w.r.t "x" ,

"\frac{dy}{dx}=3x^2+7(2)x-4(1)+0
\\\Rightarrow \frac{dy}{dx}=3x^2+14x-4" [Using "\frac{d}{dx}(x^n)=n.x^{n-1}" ]

(c):

"s=17t^3 - 5t^2 - 5t - 3"

On differentiating both sides w.r.t "t" ,

"\frac{ds}{dt}=17(3)t^2-5(2)t-5(1)-0" [Using "\frac{d}{dx}(x^n)=n.x^{n-1}" ]

"\Rightarrow \frac{ds}{dt}=51t^2-10t-5"